OpenStudy (anonymous):

What is the equation of the line that passes through(-10,8) and (5,5)?

OpenStudy (nirmalnema):

fine the slope of the line and then apply two point form...

OpenStudy (anonymous):

Ive tried everything but i keep getting B and idk why

OpenStudy (anonymous):

What am i doing wrong?

OpenStudy (akashdeepdeb):

Find their slope using rise/run. Use any of the points then. And use the point-slope of the line then you'll get your answer. Point-slope form = (y-y1) = m(x-x1) Where m is slope (x1,y1) is any of the 2 points given to you. Understood? :)

OpenStudy (nirmalnema):

slope = y2-y1 / x2-x1 and one point form is y- y1= slope(x-x1)

OpenStudy (anonymous):

Im confused...

OpenStudy (nirmalnema):

y-8 = -3/15 (x+10) what eq u will get after solving it...

OpenStudy (nirmalnema):

u will get 15y - 120 = -3x-30 divide the equation by three.. so u will get 5y - 40 = -x-10 transfer all terms to one side... you will get.. x+5y-30 which is equal to A